Two point charges of equal magnitude are 6.9 cm apart. At the midpoint of the line connecting them, their combined electric field has a magnitude of 48 N/C . Find the magnitude of the charges. q=____pC
Answer #1

Electric field, E = kq/r^2

E = 2[ kq/(r/2)^2] = 8 kq/r^2

Charge, q = Er^2/ 8k

q = (48 x 0.069^2)/(8 x 9 x 10^9) = 3.17 pC

Comment in case any doubt please rate my answer ...